### v1

# UD-English-CHILDES: A Collected Resource of Gold and Silver Universal Dependencies Trees for Child Language Interactions

Abstract
CHILDES is a widely used resource of transcribed child and child-directed speech. This paper introduces UD-English-CHILDES, the first officially released Universal Dependencies (UD) treebank derived from previously dependency-annotated CHILDES data with consistent and unified annotation guidelines. Our corpus harmonizes annotations from 11 children and their caregivers, totaling over 48k sentences. We validate existing gold-standard annotations under the UD v2 framework and provide an additional 1M silver-standard sentences, offering a consistent resource for computational and linguistic research.
UD-English-CHILDES: A Collected Resource of Gold and Silver Universal Dependencies Trees for Child Language Interactions
Xiulin Yang1  Zhuoxuan Ju1  Lanni Bu1  Zoey Liu2  Nathan Schneider1 1Georgetown University 2University of Florida {xy236, zj153, lb1437, nathan.schneider}@georgetown.edu liu.ying@ufl.edu
1 Introduction
The Child Language Data Exchange System (CHILDES) (MacWhinney, 2000) has long been a key resource for research in language acquisition, computational modeling of child language, and the evaluation of Natural Language Processing (NLP) tools. However, many analyses rely on different grammatical assumptions (e.g., Pearl and Sprouse, 2013; Szubert et al., 2024; Liu and Prud’hommeaux, 2021; Gretz et al., 2015; Sagae et al., 2007), and therefore adopt divergent annotation frameworks or standards. While most existing annotations use syntactic dependencies—in part due to the relative simplicity of annotation and parsing and the growing adoption of the Universal Dependencies (UD) framework (Nivre et al., 2016, 2020)—annotation practices remain inconsistent across datasets. This is largely due to the lack of a unified guideline for annotating children’s speech, which presents unique challenges not fully addressed by existing UD documentation.
As UD treebanks have become valuable resources in both NLP (e.g., Jumelet et al., 2025; Opitz et al., 2025) and language acquisition research (e.g., Clark et al., 2023; Hahn et al., 2020), there have been increasing efforts to parse CHILDES data using tools such as stanza (Liu and MacWhinney, 2024). However, the resulting annotation quality is often inconsistent and cannot be guaranteed. In this paper, we compile, harmonize, and manually correct major UD-style annotations of CHILDES data into a consistent, unified UD format, resulting in a gold-standard treebank of 48K sentences and 236K tokens (including, e.g., the tree in Figure 1). In addition, we construct a larger silver-standard treebank of 1M sentences and 6M tokens and report parser accuracy estimates. We publicly release both datasets.111pseudo-link-for-childes-UD
{dependency}{deptext}
[column sep=2em] And & a & a & green & one & .
[…]
Figure 1: UD tree for a child utterance from Lily (Providence corpus, sentID=16916280)
Corpus Children Gold Annotation Speakers UPOS Feats Utterances Tokens
Gold Silver Gold Silver
S+24 Adam Trees, UPOS; features from original CHILDES Adults Gold Convrtd 17,233 0 91,114 0
LP21 Eve Trees; others unspecified All Silver Silver 2,207 108,044 8,497 532,319
LP23 10 Children Trees; others unspecified All Silver Silver 34,530 1,101,061 168,284 6,461,084
UD-English-CHILDES 11 Children Trees, UPOS All Gold N/A 48,183 1,197,471 236,941 6,892,314
Table 1: Overview of CHILDES-based UD treebanks compiled in this paper. Source corpus labels (S+24, LP21, LP23) are defined in section 3. Note that there is overlap in the Adam data: S+24 figures are counts from the original dataset; for our version, these were filtered to avoid duplicates and merged with corresponding LP23 utterances.
Child Corpus Child age range Gold sents Gold toks Silver sents Silver toks
Laura Braunwald (Braunwald, 1971) 1;3-7;0 (1;3-7;0) 4,622 21,079 41,862 205,427
Adam Brown (Brown, 1973) 1;6-5;2 (1;6-5;2) 16,736 84,643 93,315 452,348
Eve Brown 1;6-5;1 (1;6-5;2) 2,207 8,497 108,044 532,319
Abe Kuczaj (Kuczaj, 1977) 2;4-5;0 (2;4-5;0) 4,167 22,437 38,630 230,489
Sarah Brown 1;6-5;2 (1;6-5;2) 5,347 23,233 104,926 517,654
Lily Providence (Demuth et al., 2006) 0;11-4;0 (0;11-4;0) 1,499 6,337 79,573 422,245
Naima Providence 1;3-3;11 (0;11-4;0) 2,534 14,360 236,350 1,422,543
Violet Providence 0;11-4;0 (0;11-4;0) 721 1,857 32,801 164,975
Thomas Thomas (Lieven et al., 2009) 2;0-4;11 (2;0-4;11) 4,240 20,333 313,550 2,039,132
Emma Weist (Weist and Zevenbergen, 2008) 2;2-4;10 (2;1-5;0) 2,423 13,730 74,825 474,460
Roman Weist 2;2-4;9 (2;1-5;0) 3,653 20,557 73,595 467,633
Table 2: Detailed statistics for each child, including counts of gold and silver annotations and their corresponding age ranges in months. Ages in the silver corpus are shown in parentheses. For source corpus URLs see Appendix A.
2 Related Work
2.1 CHILDES Corpora
CHILDES has played a central role in both language acquisition research and the development of NLP tools. In addition to specialized corpora—such as clinical datasets (Gillam and Pearson, 2004), naturalistic family interactions (Gleason, 1980), and controlled laboratory studies (Newman et al., 2016)—CHILDES supports a wide range of approaches to developmental linguistics. Many of its corpora inform foundational theories of language acquisition, particularly the poverty of the stimulus hypothesis (Chomsky, 1976). Researchers frequently use child-directed speech from CHILDES to quantify the distribution of linguistic structures that are central to these theories, such as wanna contraction (Getz, 2019), anaphoric one (Foraker et al., 2009; Pearl and Mis, 2011), auxiliary fronting (Perfors et al., 2011), and syntactic islands (Pearl and Mis, 2011). It has also been used in computational models of language acquisition (e.g., Abend et al., 2017).
CHILDES has also emerged as a valuable resource for NLP tool benchmarking and language model pretraining. Following the work of Huang (2016), studies such as Liu and Prud’hommeaux (2023) have highlighted the challenges faced by UD parsers when applied to child-directed speech, showing substantial performance gaps compared to adult data. CHILDES also supports recent research on pretraining dynamics (Feng et al., 2024) and the development of efficient language models, including in initiatives like the BabyLM Challenge (Choshen et al., 2024; Charpentier et al., 2025).
# sent_id = 17235906
# childes_toks = who's that
# corpus_name = Providence
# gold_annotation = True
# speaker_age = 37.49358303045237
# speaker_gender = male
# speaker_role = Mother
# type = question
# text = Who's that?
1-2 Who's _ _ _ _ _ _ _ _
1 Who who PRON WP _ 0 root 0:root _
2 's be AUX VBZ _ 1 cop 1:cop _
3 that that PRON DT _ 1 nsubj 1:nsubj SpaceAfter=No
4 ? ? PUNCT ? _ 1 punct 1:punct _
Figure 2: Example of a gold-annotated CoNLL-U sentence from the CHILDES-Providence corpus.
2.2 CHILDES Dependency Treebanks
Early dependency parsing research on English CHILDES data utilized a custom inventory of grammatical relations (GR; Sagae et al., 2004, 2005). These gradually evolved to address CHILDES-specific challenges (Sagae et al., 2007), and were applied to the entire English CHILDES corpus using a supervised parser (Sagae et al., 2010).
More recently, UD-style annotations have been introduced to CHILDES. Liu and Prud’hommeaux (2021) used a semi-automatic method to convert previous GR-based annotations into UD trees, focusing on child-produced speech (ages 18–27 months) from the Eve data within the Brown corpus (Brown, 1973). Subsequently, Szubert et al. (2024) developed gold-standard UD annotations by automatically transforming GR annotations and manually correcting them. Their dataset includes child-directed speech from the Adam data of the Brown corpus and the Hebrew Hagar corpus (Berman, 1990), addressing spoken-language-specific phenomena such as repetitions and non-standard vocabulary, as well as a mapping to semantics.
Building upon these efforts, Liu and Prud’hommeaux (2023) significantly expanded UD annotations to cover utterances from 10 children aged 18–66 months (Adam from the Brown corpus as well as 9 children from other corpora), incorporating both child and caregiver speech. Their work tackles complex spoken-language features, including speech repairs and restarts.
3 Annotations
This work leverages three existing UD treebanks: Szubert et al. (2024) (henceforth S+24), Liu and Prud’hommeaux (2021) (LP21), and Liu and Prud’hommeaux (2023) (LP23), summarized in Table 1. As these treebanks were already annotated, our human annotation efforts focused primarily on correcting errors and harmonizing annotations across corpora. We present post-compilation statistics in Tables 1 and 2. Table 1 summarizes the full corpus and its source contributions, and Table 2 provides per-child statistics.
3.1 Annotation Pipeline
Following Liu and Prud’hommeaux (2023), we collect CHILDES corpora using the childesr [^2] https://langcog.github.io/childes-db-website/ R package and identify whether each sentence has been previously annotated with a UD tree. Sentences with existing annotations are automatically tagged with Universal Part-Of-Speech tags (UPOS) using stanza333stanza 1.9.2 (Qi et al., 2020), while unannotated sentences are assigned both UPOS and dependency trees. Our current work focuses on correcting previously human-annotated data. To ensure conformity with UD guidelines, we run all processed sentences through the UD validation tool [^4] https://github.com/UniversalDependencies/tools/blob/master/validate.py and manually fix those that fail validation. The correction work is performed by three linguistics graduate students trained in UD annotation.
Many of the errors stem from mismatches between UPOS tags and dependency labels (as LP21 and LP23 used automatic UPOS tagging). In addition, we address format issues such as multiword tokens, spacing mismatches (e.g., SpaceAfter), and deprecated dependency relations not supported by current UD guidelines (e.g., compound:svc, obl:about_like, nmod:over_under). The 5 most common linguistic issues were as follows:
advmod tagged as ADP
This error commonly appears with phrasal verbs such as get up and take over. The original annotation assigns advmod as the dependency relation to phrasal verbs with POS tag ADP. We revise these to compound:prt, in accordance with the UD treatment of phrasal particles.
Auxiliaries tagged as VERB or PART
Auxiliaries such as be and have are frequently misclassified as main verbs or particles. In some cases, lemmas are also mislabeled—most notably, the lemma of contracted forms like ’s is incorrectly assigned as ’s rather than the appropriate auxiliary be. We correct both the POS and lemma annotations in these cases.
Lexical items tagged as PUNCT
The stanza parser often mislabels disfluent word fragments in spontaneous speech as punctuation marks (e.g., OK/INTJ Adam/PROPN ride/VERB dat/PUNCT ./PUNCT. We reassign these tokens appropriate UPOS labels based on context and speaker intent, often as interjections.
Determiner misrecognition
Ambiguous or reduced forms of determiners—such as de —are frequently misidentified as proper nouns (PROPN). We manually review these cases and reannotate them as DET when appropriate.
Function word heads with dependents
In previous treebanks, words appearing in functional relations such as case, mark, and aux have been assigned children, which violates UD’s constraint that these words should be leaf nodes. We reassign the erroneous dependents to the appropriate content heads, ensuring the structure conforms to UD’s projectivity and function word constraints.
3.2 Harmonization
As each treebank follows its own annotation guidelines, we performed a series of normalization steps to harmonize them into a consistent format. Our unified format is primarily based on LP23, with several adaptations described below.
Metadata
In our normalized CoNLL-U files, we include the following metadata fields with an example provided in Figure 2: sent_id (utterance ID retrieved via the childesr R package); childes_toks (tokenized utterance); corpus_name (original corpus name); gold_annotation (indicates whether the sentence is manually annotated); speaker_gender, speaker_role, and speaker_age (speaker metadata); text (the text aligned with the tree), and type (sentence type). Table 3 summarizes the distribution of the main sentence types and compares them with those in the UD 2.15 release of GUM (Zeldes, 2017), a multi-genre English corpus. Notably, questions occur in the CHILDES conversations at a much higher rate—they are nearly half (45%) as frequent as declarative utterances, as opposed to 9% in GUM.
Sentence Type CHILDES GUM (s_type)
declarative 31,996 7,695 (decl)
question 14,241 716 (q, wh)
imperative_emphatic 797 1,326 (imp, intj)
others 1,149 2,409
Table 3: Sentence type counts in gold CHILDES and GUM corpora. “Others” encompasses less frequent categories: trail off, interruption, self interruption, quotation next line, self interruption question, trail off question, and interruption question.
Punctuation
To bring the transcripts in line with written English conventions, we capitalize the first word of each utterance and infer sentence-final punctuation at the end of each sentence based on the sentence type provided in the metadata. [^5] The original data transcribes various kinds of prosodic information such as pauses. At present we do not retain this information or attempt to infer corresponding punctuation like commas and parentheses.
Reparandum
Each of the three treebanks defines its own subtypes for the reparandum and parataxis relations. For example, S+24 includes parataxis:repeat, while LP23 uses parataxis:discourse, labels not present in the current UD guidelines. Similarly, LP21 and LP23 annotate reparandum with subtypes such as restart and repetition to mark special utterance features of children’s speech. To ensure consistency across treebanks, we move all such subrelation information to the MISC column.
Others
As S+24 was annotated using the UD guidelines version 1.0, we convert the annotation using UD version 2.0 with a script [^6] https://github.com/UniversalDependencies/tools/tree/master/v2-conversion and manual annotation. For example, we shifted the head-dependent direction of flat in the annotations.
Metrics Children’s speech Parents’ speech Overall
LAS 81.2 86.3 84.2
UAS 87.2 91.0 89.5
Table 4: LAS and UAS scores for children’s speech, parents’ speech, and overall performance.
Since S+24 and LP23 overlap in the Adam corpus, we merged the annotations from these two treebanks. 3375 sentences are repetitive in S+24. We removed these sentences from our corpus777883 sentences from S+24 could not be merged and were therefore removed from our treebank as well..
To ensure a more linguistically plausible analysis, we also diverged from Liu and Prud’hommeaux (2023) in our treatment of interjections. Instead of annotating utterances consisting solely of interjections (e.g., Ha ha ha ha) as conj, we used the flat relation, as illustrated in Figure 3.
{dependency}
[edge unit distance=2ex] {deptext}[column sep=.5em] Ha & ha & ha & ha & .
[…]
{dependency}
[edge unit distance=2ex] {deptext}[column sep=.5em] Ha & ha & ha & ha & .
[…]
Figure 3: A tree example before (left) and after (right) standardization.
3.3 Silver Data Assessment
To create silver-standard annotations, we apply stanza to the unannotated portion of the corpus. To estimate the quality of these silver annotations, we evaluate the parser’s performance on the gold-standard data. We report Labeled and Unlabeled Attachment Scores (LAS/UAS) in Table 4. The parser achieves an overall LAS of 83.3. Performance is higher on parents’ speech (86.3 LAS) than on children’s speech (81.2 LAS), likely due to the greater syntactic regularity and lower frequency of disfluencies in adult utterances.
4 Conclusion & Future Work
In this paper, we present the first harmonized UD treebanks for CHILDES, covering 11 corpora and over 48k sentences from both child-directed and child-produced speech. Future work will involve further corrections to the silver data and the expansion of the treebanks. We invite collaboration on this ongoing effort.
Appendix A Sources of the Coprora